OpenStudy (anonymous):

\[\log_{5}x=\sqrt(\log_{5}x)\] solve the equation

OpenStudy (anonymous):

\[\log_{5}x=(\log_{5})^{1/2}\]\[\log_{5}x=1/2\log_{5}\] x=1/2 not sure,you my check some pne to answer,

OpenStudy (anonymous):

you may repost I'm not sure

OpenStudy (anonymous):

(log_5 x)^2=log_5 x, (log_5 x)^2-log_5 x=0, let u = log_5 x, u^2-u=0 implies u(u-1)=0, thus, u = 0 or 1, log_5 x =0 or log_5 x =1, x=5^0 or x=5^1.